### v1

# Process-based Indicators of Vulnerability Re-Introducing Code Changes: An Exploratory Case Study

Abstract.
Software vulnerabilities often persist or re-emerge even after being fixed, revealing the complex interplay between code evolution and socio-technical factors. While source code metrics provide useful indicators of vulnerabilities, software engineering process metrics can uncover patterns that lead to their introduction. Yet few studies have explored whether process metrics can reveal risky development activities over time — insights that are essential for anticipating and mitigating software vulnerabilities. This work highlights the critical role of process metrics along with code changes in understanding and mitigating vulnerability reintroduction. We move beyond file-level prediction and instead analyze security fixes at the commit level, focusing not only on whether a single fix introduces a vulnerability but also on the longer sequences of changes through which vulnerabilities evolve and re-emerge. Our approach emphasizes that reintroduction is rarely the result of one isolated action, but emerges from cumulative development activities and socio-technical conditions. To support this analysis, we conducted a case study on the ImageMagick project by correlating longitudinal process metrics such as bus factor, issue density, and issue spoilage with vulnerability reintroduction activities, encompassing 76 instances of reintroduced vulnerabilities. Our findings show that reintroductions often align with increased issue spoilage and fluctuating issue density, reflecting short-term inefficiencies in issue management and team responsiveness. These observations provide a foundation for broader studies that combine process and code metrics to predict risky fixes and strengthen software security.
Software security, Vulnerability reintroduction, Risky fixes, Software process metrics, Longitudinal metrics, Mining software repositories, Scientific software, Vulnerability prediction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debuggingccs: Security and privacy Software security engineering
1. Introduction
Figure 1. Overview of the vulnerability reintroduction process. Contributors submit code to an open-source project which undergoes a software review process before being committed to the version control system. After the commit is made, a vulnerability in the contribution may be identified and disclosed to a vulnerability database (e.g., CVE, NVD). After which contributors submit new code that addresses the vulnerability while simultaneously reintroducing a vulnerability. This reintroduced vulnerability is then identified where it is patched out by a contributor in a potentially cyclical pattern.
Software security remains a critical challenge within open-source software. The rapid pace and complexity of technological advancement places software engineers under time constraints, increasing the risk of flawed implementations and introducing vulnerabilities. Software vulnerabilities are reported to vulnerability databases including the Common Vulnerabilities and Exposures (CVE) (The MITRE Corporation, 2025b) and National Vulnerability Database (NVD) (National Institute Of Standards And Technology, 2025).
Despite software engineers’ efforts to address vulnerabilities, their fixes can inadvertently introduce new security issues (Śliwerski et al., 2005; Shimmi and Rahimi, 2022) as visualized in Figure˜1. Vulnerable software systems compromise overall system dependability (Braz et al., 2022), hampers reproducibility (Kaur et al., 2021), and diminishes trust (Wermke et al., 2022) Moreover, the integration of software dependencies packages enables vulnerabilities to propagate through a software supply chain (Zimmermann et al., 2019; Zahan et al., 2022; Hegewald and Beyer, 2025). Although automated security audits (Zahan et al., 2023; Snyk Limited, 2025; Google, 2025; GitHub, 2025a) and tooling (Nong et al., 2023; Alazmi and De Leon, 2022; Nie et al., 2022; Sheng et al., 2025) are available to the open-source community, the impact of these tools on engineering decisions is limited (Eisty et al., 2018; Ayala et al., 2025). While research has examined software vulnerability prediction (Nong et al., 2023; Sheng et al., 2025), a gap persists in identifying vulnerability reintroduction activities where remediation of one vulnerability inadvertently introduces another (Shimmi and Rahimi, 2022). Furthermore, as vulnerabilities are introduced and reintroduced as commit operations, a gap exists in analyzing the software engineering process involved in such operations.
To address this, we conduct the first empirical case study  of a software’s engineering process and its correlation to software vulnerability reintroduction activities. We identify pairs of vulnerability reintroducing and fixing commits and automatically compute longitudinal process metrics for bus factor, issue density, and issue spoilage for the ImageMagick (ImageMagick Studio LLC, 2024) project — a well known image processing project. Our results show that (1) ImageMagick has maintained a healthy number of contributors, implying that there is always a contributor available to fix and review vulnerability reintroducing commits; (2) ImageMagick’s issue density has remained low even during periods of vulnerability reintroduction implying that maintainers can adequately handle incoming issues while working on new features; implying that the project’s development workflow can accommodate backlog accumulation and ensuring that regular development activities are not disrupted by recurring vulnerabilities. (3) ImageMagicks’s issue spoilage typically rises during the time window of reintroduced vulnerabilities, implying that contributors are not closing issues as fast as they are opened. Periods characterized by vulnerability reintroductions coincide with increased issue spoilage, suggesting that developers may temporarily shift focus toward remediation efforts at the expense of issue closure rates. This pattern highlights a potential process bottleneck during security-intensive phases, where parallel task management may strain available resources.
Taken together, these findings suggest that while ImageMagick’s contributor base and issue management processes provide a strong foundation for sustainable maintenance, vulnerability reintroductions still impose measurable strain on project dynamics. The increase in issue spoilage during such periods indicates that even well-staffed open-source projects face coordination and prioritization challenges when balancing security remediation with ongoing development. This underscores the need for improved process-aware tooling and decision support systems that help developers anticipate, detect, and mitigate reintroduction risks before they propagate through the codebase.
Our contributions are:
  1. (1)
    The first empirical case study correlating longitudinal software engineering process metrics with open-source software engineering vulnerability reintroduction activities.
  2. (2)
    A methodology identify for identifying vulnerability reintroducing commits leveraging existing algorithms (Śliwerski et al., 2005) and LLMs.
  3. (3)
    A dataset of 76 vulnerability-reintroducing commits that extends existing vulnerability datasets by including future commits that fix the reintroduced vulnerabilities.
2. Background and Related Work
In this section, we give a brief overview software engineering process metrics, process metrics for predicting software vulnerabilities, and software vulnerability reintroduction activities.
2.1. Software Engineering Process Metrics
Software engineering process metrics quantify the activity involved in the creation or maintenance of a software system (Eisty et al., 2018). Prior work has shown that code complexity, churn, and developer activity metrics can serve as indicators of software vulnerabilities (Shin et al., 2010) however the open-source community has seen limited adoption of these and other process metrics (Eisty et al., 2018). However, these approaches primarily rely on measurements taken at specific releases, rather than examining longitudinal trends. Captured at a specific release, snapshot process metrics offer insights into the software system’s current state, however the measuring process metrics longitudinally captures socio-technical trends influencing project evolution (Synovic et al., 2022). Several process metrics have been proposed (Fritz et al., 2014; Goggins et al., 2021; Linux Foundation, [n. d.]), but for our study we focus on derived process metrics including bus factor, issue density, and issue spoilage.
Bus Factor measures team-wide familiarity with a system’s components by counting contributions to various elements (Cosentino et al., 2015); high bus factors indicate widespread understanding, while a low bus factor (i.e., less than one (Piggot and Amrit, 2013)) suggests siloed knowledge potentially limiting the number of active contributors who could resolve vulnerabilities.
Issue Density extends the defect density metric (Fenton and Bieman, 2014) to open-source issue trackers; high density suggests engineers struggle to manage community submissions efficiently, whereas low density implies timely responses to these inputs.
Issue spoilage quantifies how quickly engineering teams close reported issues; high spoilage indicates unaddressed lingering issues, whereas low spoilage signifies prompt handling of issues.
Our study advances software engineering process metrics by analyzing longitudinal process metrics, including bus factor, issue density, and issue spoilage, to correlate engineering activities vulnerability reintroduction. This approach underscores the critical gap between awareness and practical adoption of such metrics in software engineering communities, emphasizing their potential as actionable indicators for improving code quality and security resilience.
2.2. Software Vulnerabilities and Vulnerability Reintroduction
A software vulnerability refers to specific weakness in a software system’s code that can be exploited to compromise the confidentiality, integrity, or availability of a system. Centralized databases aggregate, report, score, and distribute publicly disclosed vulnerabilities to enable vulnerability tracking and awareness (National Institute Of Standards And Technology, 2025; The MITRE Corporation, 2025b; Google, 2025; GitHub, 2025b; GitLab, 2025). Vulnerabilities can be classified via the Common Weakness Enumeration (CWE) schema (The MITRE Corporation, 2025b) which captures recurring weaknesses in software (e.g., buffer overflows, improper input validation) that automated tooling can classify (Das et al., 2021; Han et al., 2017). Furthermore the Common Vulnerability Scoring System (CVSS) provides a means to quantify the severity of a vulnerability from a range of zero to ten (Forum of Incident Response and Security Teams, 2024), with ten being reserved for the most severe vulnerabilities (Benny Isaacs et al., 2025; Chen Zhaojun, 2021).
Software fixes — applied commits — attempt to address and remove the software vulnerability, but previous work has found that not all fixes are permanent (Śliwerski et al., 2005; Braz et al., 2022). Prior work has identified that vulnerability reintroduction occurs when a fix inadvertently reintroduces a new security weakness as confirmed by later fixes (Shimmi and Rahimi, 2022; Braz et al., 2022). Software vulnerability reintroduction was conceptually identified as a threat to software systems via attack–defense co-evolution by positing that a vulnerability fixing operation might itself serve as a predictor of a subsequent vulnerability (Shimmi and Rahimi, 2022). Such reintroductions pose significant risks because they may silently undo prior security improvements and leave systems exposed once again.
While existing datasets have labeled vulnerable and non-vulnerable source code (Chen et al., 2023; Fan et al., 2020), to our knowledge, there are no datasets that specifically capture vulnerability reintroducing commits and their fixes. Our work resolves this by releasing a dataset that extends established vulnerability datasets (Chen et al., 2023; Fan et al., 2020) with 81 pairs for the ImageMagick project that both resolve existing CWEs while simultaneously reintroducing new vulnerabilities.
3. Case Study: ImageMagick
In this section we introduce our case study on ImageMagick — a well known and embeddable image processing project. We discuss an overview of the problem, our methodology to evaluate the project’s software engineering process metrics and high CVSS scoring CVEs longitudinally, and analyze the data collected during our case study. We selected ImageMagick as our case study due to it being a well known open-source image processing project with a significant development history (ImageMagick Studio LLC, 2024), written primarily in C potentially enabling common weaknesses (e.g., buffer overflows, memory leaks), and being readily available in existing vulnerability datasets that have previously mapped CVE disclosures to specific commits (Chen et al., 2023; Fan et al., 2020).
3.1. Problem Overview
In this section we show an example of a vulnerability that was reintroduced in the fix for a separate vulnerability. CVE-2018-11625 (The MITRE Corporation, 2018a) was an out of bounds read vulnerability (The MITRE Corporation, 2025c) with a CVSS severity score of 8.8 disclosed on May 31st, 2018.
Vulnerability Reintroducing Fix
Listing 1 shows a commit accepted by ImageMagick that reintroduced a vulnerability while attempting to resolve this vulnerability. On May 30th, 2018 commit 5294966 was accepted into ImageMagick to resolve a buffer overflow by allocating an additional byte for colormap_index. While the intention was to prevent out-of-bounds writes, this fix was potentially unsafe when image->colors or MaxColormapSize was smaller than MaxMap, creating scenarios where the buffer could still overflow under certain image configurations. [^1] https://github.com/ImageMagick/ImageMagick/commit/5294966
Listing 1: ImageMagick – Initial Fix Commit 5294966 - colormap_index=(unsigned short *) AcquireQuantumMemory( - (size_t) image->colors, sizeof(unsigned short)); + colormap_index=(unsigned short *) AcquireQuantumMemory( + (size_t) (image->colors+1), sizeof(unsigned short));
Reintroduced Vulnerability Fix
Listing 2 shows the subsequent commit c111ed9 that corrected the reintroduced vulnerability on April 8th, 2019. This issue was resolved by ensuring that the allocation size was properly bounded using the MagickMax() function, guaranteeing that the memory allocated for colormap_index was at least MaxMap. This change prevented buffer overflow conditions that could occur when the color count was smaller than the maximum mapping size. [^2] https://github.com/ImageMagick/ImageMagick/commit/280215b9936d145dd5ee91403738ccce1333cab1
The reintroduced vulnerability remained active for 313 days across 1,094 commits and 6 releases.
Listing 2: ImageMagick – Future Fix Commit c111ed9 - colormap_index=(unsigned short *) AcquireQuantumMemory( - (size_t) (image->colors+1), sizeof(unsigned short)); + colormap_index=(unsigned short *) AcquireQuantumMemory( + MagickMax((size_t) image->colors, MaxMap), sizeof(unsigned short));
Figure 2. Our methodology to extract and compute longitudinal SEP metrics for vulnerability reintroducing and fixing commits in the ImageMagick project. We extract issues and pull requests from the GitHub issue tracker for ImageMagick to compute issue density and issue spoilage. Git commits are leveraged to compute bus factor and issue density, and to identify vulnerability fixing commits with the SZZ algorithm and Anthropic Claude 3.5 Sonnet from the ImageMagick GitHub repository, BigVul, and DiverseVul. Data is also collected from CVE and is leveraged to visualize the data.
3.2. Methodology
To evaluate the socio-technical context in which vulnerabilities are reintroduced into the ImageMagick dataset we extract and evaluate vulnerability resolving commits from existing datasets, extend existing algorithms and methods with LLMs to identify pairs of reintroducing and fixing commits, and leverage prior work on evaluating longitudinal software engineering process metrics on selected high CVSS scoring CVEs that were identified to reintroduce issues. Our goal is twofold: (1) to flag vulnerability fixing commits that are likely to trigger regressions, and (2) to isolate the engineering process patterns that contribute to vulnerability reintroducing activity in open-source software. Our methodology is outlined in Figure˜2
3.2.1. Dataset Collection
As current vulnerability datasets do not provide vulnerability reintroducing and fixing commit pairs, we developed a ground truth dataset for ImageMagick that extends existing vulnerability datasets with vulnerability reintroducing commits. We first extracted existing ImageMagick vulnerability commits from BigVul (Fan et al., 2020) and DiverseVul (Chen et al., 2023), two existing and well-studied vulnerability datasets. For our initial case study, we filtered for commits that modified only one file. To filter for commits that specifically reintroduced vulnerabilities, we analyzed subsequent commits to determine whether regressions occurred that modified the initial fix with the SZZ algorithm (Śliwerski et al., 2005). As SZZ solely operates on source code, we also manually evaluated the commit message attached to SZZ identified commits. To further refine our dataset we leveraged Anthropic Claude 3.5 Sonnet (https://claude.ai/) to confirm if each vulnerability reintroducing and fixing commit was accurate. Our prompt to do so is presented in Listing 3.
Listing 3: Anthropic Claude 3.5 Sonnet System Prompt Research context: I am conducting research on how software security vulnerabilities evolve over time, specifically focusing on situations where a fix for one security vulnerability unintentionally introduces a new security vulnerability. Task: You will be provided with details from two commits: Previous Fix Commit – A commit that fixed a known vulnerability. Future Candidate Commit – A later commit that modifies the same or nearby code. Your task is to determine whether the candidate commit is fixing a new vulnerability that was introduced by the previous fix. Previous Fix Details Commit ID: {commit_hash} Commit Message: {previous_fix_message} Code Changes (Diff Format): {previous_fix_diff_content} Future Candidate Details Commit Message: {future_commit_message} Code Changes (Diff Format): {future_diff_content} Your Response Format (Strictly Follow This JSON Format) {"answer": "Yes" or "No",
"reasoning": "Detailed explanation of why the candidate commit is or is not fixing a vulnerability introduced by the previous fix."
Important condition: If the previous fix is incomplete, the answer is "No" since it did not introduce a new vulnerability. Similarly, if the previous fix did not properly fix the issue, the answer is still "No" unless a new vulnerability is created.
Figure 3. Anthropic Claude 3.5 Sonnet LLM system prompt used to evaluate commit pairs where one commit fixed a vulnerability and a subsequent commit potentially addressed a new vulnerability introduced by that fix.
Our dataset collection resulted in 81 confirmed vulnerability reintroducing-fixing commit pairs for ImageMagick from an initial set of 175 and 225 commits from BigVul and DiverseVul respectively.Of the 81 pairs, 25 originated from BigVul (Fan et al., 2020) and 56 originated from DiverseVul (Chen et al., 2023), however only 76 have associated CVEs. Our dataset is publicly available [^3] https://github.com/anonSubmissionGithub/ProcesseMetric-VulReintroduction.
3.2.2. CVE Selection Criteria
Table 1. CVEs analyzed and their CWE vulnerability classification and CVSS severity score.
CVE CWE CVSS
CVE-2016-4564 CWE-119 9.8
CVE-2017-16546 CWE-119 8.8
CVE-2018-11625 CWE-125 8.8
CVE-2019-13299 CWE-125 8.8
From the 76 candidate pairs of vulnerability reintroducing commits, we randomly selected four pairs. Our selection criteria were that each pair was identified to resolve a known CVE with at least a score of 8.8 while reintroducing a new vulnerability, that each vulnerability pair is non-overlapping, and each vulnerability had to be identified in a unique year. Our candidate vulnerabilities are presented in Table˜1.
3.2.3. Software Process Metrics Analysis
For longitudinal computation of process metrics, we utilized the PRIME tool (Synovic et al., 2022). PRIME can compute bus factor, issue density, and issue spoilage among others for any GitHub hosted source code repository with a GitHub issue tracker. For bus factor, we computed the bus factor for the entire project every six months. For issue density and spoilage, we computed the metric per six months, but for each CVE in our case study we computed the metric per week for 20 weeks prior to, during, and 20 after the reintroduced vulnerability persisted in the project.
3.3. Data Analysis
In this section we discuss an overview of the dataset of 76 candidate vulnerability reintroducing-fixing commit pairs identified in Section˜3.2.1, and our longitudinal analysis of bus factor, issue density, and issue spoilage as it relates to vulnerability reintroduction activities.
3.3.1. ImageMagick Vulnerabilities per Year:
Table 2. Breakdown of the number of vulnerability reintroducing commits with identifiable CVEs in ImageMagick from 2015 to 2021.
Year # Of Vuln. Commits Avg. CVSS
2015 5 6.5
2016 12 7.4
2017 24 7.1
2018 10 6.8
2019 11 7.7
2020 13 5.0
2021 1 5.5
Table˜2 breaks down the 76 out of 81 identified vulnerability reintroducing commits with CVEs and their average CVSS scores per year. Since 2017, the number of vulnerabilities that are reintroduced into ImageMagick has decreased. Furthermore, the average CVSS score of vulnerabilities that reintroduce vulnerabilities has also decreased. This trend suggests that ImageMagick’s contributors are becoming increasingly effective at identifying, isolating, and mitigating security flaws before they escalate in severity. It also indicates a maturation of the project’s secure development practices, where vulnerability resolution attempts are less likely to result in the reintroduction of high-impact security issues.
3.4. Bus Factor
Figure 4. ImageMagick maintains a relatively healthy bus factor, with contributions from 4-7 unique maintainers every six months, facilitating distributed handling of vulnerability resolutions.
ImageMagick’s bus factor per six months metric is presented in Figure˜4. ImageMagick’s has had at least four contributors per six months with as many seven contributing to the project. This guarantees that at any one time, there have been a sufficient number of contributors who can resolve vulnerabilities. It also implies that there is a probability that any any given contributor could reintroduce a vulnerability while attempting to resolve a CVE.
3.5. Issue Density
Figure 5. ImageMagick’s issue density peaked in 2022 with 28% issue density, but has since dropped to 16.4% as of writing. CVE-2016-4564 saw increasing issue density because the rate of new issues being opened exceeded the maintainers’ closure rate. CVE-2017-16546 and CVE-2018-11625 both experienced density dips due to code deletion and the closing of numerous issues, respectively. CVE-2019-13299 showed increasing density driven by a rise in open issues.
ImageMagick’s and selected CVEs’ issue density is presented in Figure˜5. ImageMagick’s issue density per six months is low, at only 28% of open issues to total project size at its peak in 2020.
CVE-2016-4564: Prior to the vulnerability being reintroduced, issue density was increasing. After the vulnerability was reintroduced, issue density slightly decreased. After the reintroduced vulnerability was fixed, issue density decreased weekly, until rising once more. When evaluating the rate at which the project size was changed — measured in thousands of lines of code per week (i.e., KLOC per week) — during this time frame, we see that it experiences little change. Thus, the issue density of this CVE is reflective of the contributors initially closing issues raised by the community, before the community started contributing more issues than the author’s could close.
CVE-2017-16546: Prior to the vulnerability being reintroduced, issue density was decreasing. After the vulnerability was reintroduced, issue density experienced a wave of increasing issue density, before issue density decreased. After the reintroduced vulnerability was fixed, issue density increased. When evaluating the KLOC per week during this time frame, we see that it dips early in the time frame before returning to normal. Thus, the issue density of this CVE is reflective of code being removed from the project before being added back, resulting in a noticeable spike in issue density at this time.
CVE-2018-11625: Prior to the vulnerability being reintroduced, issue density was increasing. After the vulnerability was reintroduced, issue density continued to rise for a five weeks before sharply decreasing. This followed by another increase in issue density, followed by a decrease, then continuous increasing throughout the remainder of the time frame and after the fix. When evaluating the KLOC per week during this time frame, we see that it increases during this time frame. Thus, the issue density of this CVE is reflective of the project maintainers both modifiying the code and closing issues concurrently during this time frame.
CVE-2019-13299: Prior to and during the vulnerability reintroduction, and after the fix was released, issue density was increasing. When evaluating the KLOC per week during this time frame, we see that it is stable. Thus, the issue density of this CVE is reflective of an increasing amount of issues being opened during this time frame.
Issue density exhibited three distinct trajectories across our four selected CVEs:
  • Declining stability (1/4 CVEs – CVE-2016-4564): Density decreased after the risky fix, implying improved triage and faster issue closure.
  • Fluctuating recovery (2/4 CVEs – CVE-2017-16546, CVE-2018-11625): Density oscillated due to simultaneous code refactoring and issue handling.
  • Persistent growth (1/4 CVEs – CVE-2019-13299): Density continuously increased, indicating sustained issue accumulation and limited closure capacity.
These findings suggest that vulnerability reintroduction is often accompanied by temporary disruptions in issue management efficiency, reflecting the project’s fluctuating capacity to balance bug triage and ongoing development. This observation is particularly interesting and warrants further investigation across larger and more diverse projects.
3.6. Issue Spoilage
Figure 6. ImageMagick’s issue spoilage peaked in 2023 with 74,657 days of spoilage, but has since dropped to 25 days as of writing. CVE-2016-4564 saw an increase in issue spoilage after the vulnerability reintroduction, dipping only after the corrective fix was published. CVE-2017-16546 and CVE-2018-116 both showed a dip in issue spoilage immediately after vulnerability reintroduction. CVE-2019-13299 experienced increasing issue spoilage, implying that ImageMagick was unable to keep up with closing related issues in a timely manner.
ImageMagick’s and selected CVEs’ issue spoilage is presented in Figure˜6. ImageMagick’s issue spoilage per six months is high but showing improvement, with over 74,657 days of spoiled issues at its peak in 2023, and currently at 25 days of spoiled issues as of writing.
CVE-2016-4564: Prior to and during the reintroduced vulnerabilities time frame, issue spoilage was increasing. After the reintroduced vulnerability was fixed, issue spoilage decreased before before increasing again. This reflects that towards the tail end of the reintroduction period, the project maintainers put in the effort to resolve long open issues before resuming other activities.
CVE-2017-16546: Prior to the vulnerability being reintroduced, issue spoilage was decreasing. After the vulnerability was reintroduced, issue spoilage experienced a wave of increasing issue spoilage, a dip, and then increasing issue spoilage. After the reintroduced vulnerability was fixed, issue spoilage continued to increase. This reflects that during this reintroduced vulnerabilities life, the maintainers were actively resolving issues while conducting other activities.
CVE-2018-11625: Prior to the vulnerability being reintroduced and initially after the introduction, issue spoilage was increasing. During the reintroduction time frame, issue spoilage experienced a wave of increasing issue spoilage, a dip, and then continuous increase throughout the time frame and after. This shows that the maintainers initially closed a significant number of issues, and then continued to keep up with resolving open issues throughout the vulnerabilities reintroduction period.
CVE-2019-13299: Prior to and during the vulnerability reintroduction, and after the fix issue spoilage was increasing. This is reflective of the maintainers not being able to close open issues faster than they were being opened.
Issue spoilage exhibited three distinct patterns across our four selected CVEs:
  • Reactive escalation (1/4 CVEs): Spoilage increased after risky fix
  • Aggressive velocity (2/4 CVEs): Spoilage declined after risky fix
  • Sustained degradation (1/4 CVEs): Spoilage continuously increased
These patterns collectively indicate that fluctuations in issue spoilage correspond to varying levels of responsiveness and process stability during vulnerability reintroduction periods.
4. Threats To Validity
We report our findings based on an initial study of a subset of vulnerability reintroduction commit pairs from ImageMagick — a well known and embeddedable image processing project. The relatively small sample size of analyzed commti pairs limits the generalizability of our conclusions. Since ImageMagick is primarily written in C/C++, the results may not generalize to projects in other languages or domains. Furthermore, not all open-source projects publicly disclose vulnerabilities to centralized vulnerability databases, potentially limiting the applicability of our approach to the wider open-source community. To mitigate these threats we relied on existing vulnerability datasets (Fan et al., 2020; Chen et al., 2023) to identify projects with identified CVE and CWEs to focus on projects of interest to the software engineering community.
Our methodology to identify vulnerability reintroducing and fixing commit pairs leveraged Anthropic Claude 3.5 Sonnet LLM to confirm manually identified commit pairs from the BigVul vulnerability dataset (Fan et al., 2020). We also leveraged the LLM to autonomously identify commit pairs from the DiverseVul vulnerability dataset (Chen et al., 2023). This decision was driven by the significant time and complexity involved in conducting a full manual analysis.
Furthermore, our leveraged bus factor metric (Cosentino et al., 2015) measures the aggregate number of unique contributors to a project, and does not take into account who contributes to individual directories or files. Thus, while our findings show that at least four contributors are working on ImageMagick at any one time, it does not enable the evaluation of the engineering activity of each contributor. To mitigate this, we do not focus on the engineering contributions of any one contributor, but rather looking at the wholistic effort of the contributing team.
5. Discussion and Future Implications
This initial case study on ImageMagick demonstrates the value of analyzing SEP metrics to understand vulnerability reintroduction patterns. Our findings reveal potential insight on how bus factor, issue density, and issue spoilage correlate with vulnerability reintroduction activities. These findings highlight several directions for future research.
Expanding the Scope Beyond ImageMagick First, we will expand the scope of analysis to multiple open-source systems across diverse domains and programming languages. By extending our methodology to diverse projects, we want to assess whether there is any general correlation between SEP and vulnerability reintroduction. This large-scale analysis will allow us to determine if certain socio-technical configurations are consistent among projects.
Developing Predictive Models for Risky Fixes The overarching objective of this research is to enable actionable prediction of vulnerability reintroduction. By integrating longitudinal process metrics with code-level features and commit metadata, we aim to develop machine learning classifiers capable of identifying vulnerability-fixing commits that exhibit a high likelihood of reintroducing new security issues at the time they are proposed. Such predictive systems could inform automated regression testing and security-focused quality assurance workflows.
Investigating Pull Request Dynamics Future work will examine how code review processes influence vulnerability reintroduction. We will analyze metrics such as pull request spoilage, review latency, reviewer-to-contributor ratios, and comment volume to assess their correlation with reintroduction risk. Understanding these dynamics will help determine whether certain review practices, such as rushed approvals, limited reviewer diversity, or insufficient review iterations, are associated with higher vulnerability reintroductions.
6. Conclusion
This work aims to highlight the critical role of software engineering process metrics in understanding and mitigating the risk of vulnerability reintroduction. By shifting the focus from file-level vulnerability prediction to commit-level analysis of security fixes, we underscore the importance of detecting risky changes not only at the moment they are introduced, but also across long sequence of changes that capture how vulnerabilities evolve and re-emerge over time. Our approach emphasizes that vulnerability reintroduction is rarely the result of a single isolated action, but instead emerges from cumulative sequences of development activities and socio-technical conditions.